help on
One numbers is 2 more than 2 times the other, and their sum is 27. what are the numbers.

Answer by cathieb1225(41)

One numbers is 2 more than 2 times the other, and their sum is 27. what are the numbers.
x is the first number
2x+2 is the second number
x+2x+2=27
Add like terms
3x+2=27
Now subtract 2 from each side
3x+2-2=27-2
3x=25
Now divide both sides by 3
3x/3=25/3
x=25/3 or 8 1/3 (first number)
2x+2=18 2/3 (second number)